OpenStudy (anonymous):

how we can equate (2)^1/2 sinthea from 1/2(costheta)

OpenStudy (anonymous):

sqrt 2 sin theta = 0.5 cos theta? sqrt 2 tan theta = 0.5 tan theta = 1 / (2 sqrt 2) = (2 sqrt 2) / (2 sqrt 2)^2 = 2 sqrt 2 / 8 = sqrt 2 / 4